A local electrical appliances shop has found from the experience that the demand of the tube lights is distributed as Poison with mean of 4 tube lights per week. If the shop keep 6 tubes during a particular week, what is the probability that the demand will exceed the supply during that week.


1
Expert's answer
2021-03-10T11:28:55-0500

Let "X=" the demand of the tube lights: "X\sim Po(\lambda)."


"P(X=x)=\dfrac{e^{-\lambda}\cdot\lambda^x}{x!}"

Given "\lambda=4"


"P(X>6)=1-P(X\leq6)"

"=1-0.88933=0.11067"

The probability that the demand will exceed the supply during that week is 0.11067.